Your organization is planning to purchase a new office building five years from now. The building will cost $3,000,000, have a useful life of 25 years, and have no salvage value. If your organization puts aside $550,000 in an investment account with an annual interest rate of 3.75%, how much additional money must your organization deposit into the account at the end of each month in order to be able to purchase the building in exactly five years?
Answer #1

Future value of $550,000 set aside = $550,000 * (1+3.75%) 5 = 661155 [ rounded off]

Balance future value required = 3,000,000 - 661155= 2338845

Formula to find PMT :

FV of annuity = PMT *[ (1+r) n -1] /r

2338845 = PMT *[ (1+3.75%/12) 60 -1] /(3.75%/12)

PMT = 35501.14

deposit into the account at the end of each month in order to be able to purchase the building in exactly five years = $35501.14